# IOPO: Empowering LLMs with Complex Instruction Following via Input-Output Preference Optimization

### v2

@@ 6.2 Experimental Results
Main Results. As shown in Table 3, we give the main results under different benchmarks, including in-domain Trace, out-of-domain IFEval and CFBench. The experiments are conducted under two different base models, Qwen2-7B, and Llama3.1-8B, where Instruct means directly using Qwen2-7B-Instruct or Llama3.1-8B-Instruct for inference, SFT represents the model is trained on Trace training set, and PPO, DPO, IOPO are respectively trained on preference data derived from Trace training set.
For in-domain evaluation on Trace set, we can see 3.0%, 1.7% improvements of IOPO on single- and multi-constraint instructions with Qwen2-7B as the base model compared to DPO, and 2.5%, 1.5% improvements with Llama3.1-8B as the base model. For out-of-domain evaluation on IFEval and CFBench, IOPO achieves an average increase of 3.2%, and 3.06% in comparison with DPO based on Qwen2-7B and Llama3.1-8B respectively. The significant advantages of both in-domain and out-of-domain evaluations confirm the effectiveness of input-output preference optimization, which intensively considers the constraint differences between instructions, enhancing the model’s perception of constraints. It is worth noting that IOPO has a larger performance gap with SFT especially on IFEval, compared to DPO and SFT, which confirms the generalization of IOPO and the necessity of further modeling input preferences.
Ablation Studies. To further confirm the effectiveness of input and output preference, we conduct the ablation studies on Trace, IFEval, and CFBench as shown in Table 4, where “w/o Output Pref” means we only consider the modeling of input preference with the same training data, “w/o Input Pref” means we only consider the modeling of output preference. We can see that output preference contributes to 2%, and 0.93% increases with Qwen2-7B and Llama3.1-8B respectively, input preference separately brings 1.51% and 1.58% performance gains, which confirms the effectiveness of both input and output preference modeling. Besides the paradigm for modeling output preference in existing alignment methods, it’s established that modeling input preference is crucial for deeply considering constraints within the instruction.
@@
Complexity Analysis. We conduct the analyses of complexity in Table 5, where all methods are conducted under the same experimental settings, such as the batch size and GPU. (1) For #Memory, DPO and IOPO are approximately twice and four times that of SFT respectively, because DPO needs a pair of responses to calculate the corresponding loss (<$x,y_{1}$>, <$x,y_{2}$>), and IOPO needs to compute four groups of input-output pairs (<$x_{1},y_{1}$>, <$x_{2},y_{2}$>, <$x_{1},y_{2}$>, <$x_{2},y_{1}$>) in its loss. (2) For #Training Time, DPO and IOPO require the computation of more tokens compared to SFT under the same batch size, leading to longer training time. (3) For #Inference Speed, SFT, DPO, and IOPO are all the same base model optimized for inference, resulting the same inference speed. The training efficiency and GPU memory usage of IOPO are not the best among compared baselines, but their efficiencies are still of the same order of magnitude, which are reasonable and acceptable comprehensively considering the significant performance advantage.
@@
The Impact of Token Quantity. To address concerns regarding the IOPO training tokens, we conduct the analyses on the impact of token quantity and report the results in Figure 4, and Figure 5. For IOPO, there exist two instructions along with their corresponding responses ({<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}). To ensure that DPO and IOPO consume the same number of tokens, we construct two pairs of output preferences based on IOPO’s instructions $x_{1}$ ($y_{\rm win}=y_{1}$, $y_{\rm loose}=y_{2}$), and $x_{2}$ ($y_{\rm win}=y_{2}$, $y_{\rm loose}=y_{1}$) for training DPO model, denoted by DPO. Similarly, we train SFT model with instruction data {<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}, denoted by SFT. We can observe that increasing the token quantity does indeed yield better performance on some datasets. For example, compared to DPO, DPO has achieved a performance improvement on IFEval (S-Acc, L-Acc) with Qwen2-7B as the base model. However, there are also cases of decline, such as comparing DPO with DPO on CFBench (CSR, PSR), which indicates that it is not the case that more tokens always lead to better performance. At the same time, although consuming the same number of tokens, SFT and DPO still have certain gaps compared to the proposed IOPO, which confirms that the performance improvement of our IOPO does not primarily come from using more tokens, but rather from better constraint-aware modeling of input-output preferences.
@@ 7 Conclusion
In Trace, the evaluation set has undergone strict manual verification but the training set has not performed this process considering the cost. Although the models trained on the training set have achieved the significant improvements, we believe that if we can further improve the quality of the training set, it will lead to better model performance.
@@ Appendix B Prompt
Constraint Expansion Prompt:
/* Task Prompt */
@@
<explain>:
Modify the original instruction by adding constraints from two types: text style and numerical.
<result>:
Introduce Beijing to me, requiring a concise summary with no more than 1000 words, and bold any numbers and place names mentioned in the content.
—INPUT—
<instruction>:
Provide a short article about species extinction that includes at least 3 causes and 3 solutions.
—OUTPUT—
<explain>:
Modify the original instruction by adding constraints from three types: numerical, citation and reference, and inclusion.
<result>:
Provide a short article on species extinction. The length of the article should be between 500 and 800 words, detailing at least 3 reasons for species extinction and proposing at least 1 specific solution for each reason. The article should include an introduction and conclusion, and cite at least 2 scientific papers as support in the discussion.
/* Input */
—INPUT—
<instruction>:
—OUTPUT—
Instruction Structuring Prompt:
/* Task Prompt */
@@
Help me answer by introducing Beijing, where the format of the generated answer needs to be JSON, the character length of the data field in the JSON object must not exceed 1000 characters to ensure the answer is concise and direct; the generated answer must be objective facts and cannot include any subjective opinions or conjectures; all characters must be UTF-8 encoded to ensure support for multilingual responses; the answer must not contain any personal privacy information, copyright-restricted content, or information that violates relevant laws and policies; numbers, place names mentioned in the answer need to be specially marked in bold.
@@
#Task Description:
Introduce Beijing
#Constraints:
The format of the generated answer needs to be JSON;
The answer style is concise, with a character length not exceeding 1000 characters;
All characters must be UTF-8 encoded to ensure support for multilingual responses;
The generated answer must be objective facts and cannot include any subjective opinions or conjectures;
The answer must not contain any personal privacy information, copyright-restricted content, or information that violates relevant laws and policies;
Numbers, place names mentioned in the answer need to be specially marked in bold;
#Input:
NULL
/* Input */
@@
Judge Completeness Prompt:
/* Task Prompt */
/* Requirements */
The given instruction consists of three parts: <Task Description, Constraints, Input>, where Input can be NULL; You can refer to the examples given in Example, but you should not directly copy the examples;
/* Example */
@@
#Task Description:
Select the correct word from multiple options to fill in the blank that fits the context and sentence flow, and briefly explain the reason for the choice.
#Constraints:
The chosen word must fit the overall logic of the sentence;
The selected word must not only be grammatically correct but also fit the context of the sentence;
The reason for the choice must be briefly explained in parentheses;
#Input:
My car is _____ and blue.
@@
<explain>:
The Task Description mention selecting the correct word from multiple choices to fit the context and sentence flow, but the instruction does not provide any options, so this instruction is incomplete.
<tag>:
Incomplete
—INPUT—
#Task Description:
Calculate the area of a US dollar bill.
#Constraints:
The calculation result should be rounded to two decimal places;
The answer should detail each step of the calculation process.
#Input:
The length of a US dollar bill is 6.14 inches, and the width is 2.61 inches.
—OUTPUT—
<explain>:
The Task Description mention calculating the area of a US dollar bill, and the Inputs provide the length of the US dollar bill as 6.14 inches and the width as 2.61 inches. Additionally, the Constraints section provides several conditions, so this instruction is complete.
<tag>:
Complete
/* Input */
—INPUT—
—OUTPUT—
Judge Redundancy Prompt:
/* Task Prompt */
You are the redundancy detector for instructions, capable of determining whether given instructions are redundant;
/* Requirements */
The given instructions consist of <Task Description, Constraints, Input>, where Input can be NULL; You can refer to the Examples provided, but you should not directly copy the examples;
/* Example */
—INPUT—
#Task Description:
Provide a detailed explanation of the phrase “on an even keel”, including its origin and modern usage in English, presented in a professional academic tone and formatted in Markdown.
#Constraints:
The explanation must be given in a professional academic tone;
The output format must be Markdown;
The explanation must include the origin of the phrase and its usage in modern English;
#Input:
NULL
—OUTPUT—
<explain>:
The Task Description mention “including its origin and modern usage in English, presented in a professional academic tone and formatted in Markdown”, which overlaps with Constraints 1, 2, and 3. Therefore, this instruction is redundant.
<tag>:
Redundant
—INPUT—
#Task Description:
Introduce a new product line to consumers who have a high interest in environmental protection, emphasizing the environmental characteristics of the products and how they enhance quality of life.
#Constraints:
The introduction should be enthusiastic and inspiring;
The article length should be controlled between 200 and 300 words to maintain reader attention;
Clearly state the core selling points: “causing no harm to the environment and influencing your life”;
Elaborate on the specific meanings of “the revolution of sustainable living” and “being the change you wish to see”.
#Input:
NULL
—OUTPUT—
<explain>:
There is no repetition between the information mentioned in the Task Description and the four constraints listed under Constraints, so this instruction is not redundant.
<tag>:
Not Redundant
/* Input */
—INPUT—
—OUTPUT—
Response Generation Prompt:
{Instruction}
Response Evaluation Prompt:
@@
5. It is necessary to strictly follow the format in the /* Example */ for generation, the Fine Grained Score format is Json, and Constraints Overall Score format is List.
@@
/* Example */
—INPUT—
#Task Description:
Create a password for this account
#Constraints:
The password must be at least 8 characters long;
It must contain 1 uppercase letter;
It must contain 1 lowercase letter;
It must include 2 numbers;
#Input:
NULL
#Response:
Ax7y4gTf
—OUTPUT—
Explanation:
Password Length: The password “Ax7y4gTf” is 8 characters long, meeting the first constraint, scoring 10 points.
Contains 1 uppercase letter: The password “Ax7y4gTf” contains two uppercase letters, “A” and “T”, which means it meets the second constraint, but the explanation incorrectly states it does not meet the constraint, scoring 0 points.
Contains 1 lowercase letter: The password “Ax7y4gTf” contains three lowercase letters, “x”, “y”, and “g”, which means it meets the third constraint, but the explanation incorrectly states it does not meet the constraint, scoring 0 points.
Includes 2 numbers: The password “Ax7y4gTf” includes two numbers, “7” and “4”, meeting the fourth constraint, scoring 10 points.
Fine Grained Score: [ { "The password must be at least 8 characters long": 10, "It must contain 1 uppercase letter": 0, "It must contain 1 lowercase letter": 0, "It must include 2 numbers": 10 } ]
Constraints Overall Score: [[5]]
/* Input */
@@ Appendix D IOPO Data Construction
Appendix E Derivation for $p(\mathcal{G}_{1}\succ\mathcal{G}_{2})$

### v3

@@ 6.2 Experimental Results
Main Results. As shown in Table 3, we give the main results under different benchmarks, including in-domain Trace, out-of-domain IFEval, CFBench and ComplexBench. The experiments are conducted under two different base models, Qwen2-7B, and Llama3.1-8B, where Instruct means directly using Qwen2-7B-Instruct or Llama3.1-8B-Instruct for inference, SFT represents the model is trained on Trace training set, and PPO, DPO, IOPO are respectively trained on preference data derived from Trace training set.
For in-domain evaluation on Trace set, we can see 3.0%, 1.7% improvements of IOPO on single- and multi-constraint instructions with Qwen2-7B as the base model compared to DPO, and 2.5%, 1.5% improvements with Llama3.1-8B as the base model. For out-of-domain evaluation on IFEval, CFBench and ComplexBench, IOPO achieves an average increase of 2.95%, and 2.72% in comparison with DPO based on Qwen2-7B and Llama3.1-8B respectively. The significant advantages of both in-domain and out-of-domain evaluations confirm the effectiveness of input-output preference optimization, which intensively considers the constraint differences between instructions, enhancing the model’s perception of constraints. It is worth noting that IOPO has a larger performance gap with SFT especially on IFEval, compared to DPO and SFT, which confirms the generalization of IOPO and the necessity of further modeling the input preferences.
Ablation Studies. To further confirm the effectiveness of input and output preference, we conduct ablation studies on Trace, IFEval, CFBench, and ComplexBench as shown in Table 4, where “w/o Output Pref” 444 $\mathcal{L}_{\text{IOPO}^{*}}(\pi_{\theta})=-\mathbb{E}_{i\sim D}{\log[\sigma({\Pi^{*}(\pi_{\theta})})]}$ $i=<x_{1},y_{1},x_{2},y_{2}>$ $\Pi^{*}(\pi_{\theta})=\beta\log\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\text{ref}}(y_{1}|x_{1})}-\beta\log\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\text{ref}}(y_{1}|x_{2})}$ where the preference pair data $y_{2}|x_{2}$, $y_{2}|x_{1}$ are also used for training. means we only consider the modeling of input preference with the same training data, “w/o Input Pref”  [^5] Different from “w/o Output Pref”, $\Pi^{*}(\pi_{\theta})=\beta\log\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\text{ref}}(y_{1}|x_{1})}-\beta\log\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\text{ref}}(y_{2}|x_{1})}$ where the preference pair data $y_{2}|x_{2}$, $y_{1}|x_{2}$ are also used for training. means we only consider the modeling of output preference. We see that output preference contributes to 2.1%, and 1.28% increases with Qwen2-7B and Llama3.1-8B respectively, input preference separately brings 1.5% and 1.4% performance gains, which confirms the effectiveness of both input and output preference modeling. Besides the paradigm for modeling output preference in existing alignment methods, it’s established that modeling input preference is crucial for deeply considering constraints within the instruction.
@@
Complexity Analysis. We conduct the analyses of complexity in Table 5, where all methods are conducted under the same experimental settings, such as the batch size and GPU. (1) For #Memory, DPO and IOPO are approximately twice and four times that of SFT respectively, because DPO needs a pair of responses to calculate the corresponding loss (<$x,y_{1}$>, <$x,y_{2}$>), and IOPO needs to compute four groups of input-output pairs (<$x_{1},y_{1}$>, <$x_{2},y_{2}$>, <$x_{1},y_{2}$>, <$x_{2},y_{1}$>) in its loss. (2) For #Training Time, DPO and IOPO require the computation of more tokens compared to SFT under the same batch size, leading to longer training time. (3) For #Inference Speed, SFT, DPO, and IOPO are all the same base model optimized for inference, resulting the same inference speed. The training efficiency and GPU memory usage of IOPO are not the best among compared baselines, but their efficiencies are still of the same order of magnitude, which are reasonable and acceptable comprehensively considering its performance advantage.
@@
The Impact of Token Quantity. To address concerns regarding the IOPO training tokens, we conduct the analyses on the impact of token quantity and report the results in Figure 4, and Figure 5. For IOPO, there exist two instructions along with their corresponding/right responses ({<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}). To ensure that DPO and IOPO consume the same number of tokens, we construct two pairs of output preferences based on IOPO’s instructions $x_{1}$ ($y_{\rm win}=y_{1}$, $y_{\rm lose}=y_{2}$), and $x_{2}$ ($y_{\rm win}=y_{2}$, $y_{\rm lose}=y_{1}$) for training DPO model, denoted by DPO. Similarly, we train SFT model with instruction data {<$x_{1}$, $y_{1}$>, <$x_{2}$, $y_{2}$>}, denoted by SFT. We can observe that increasing the token quantity does indeed yield better performance on some datasets. For example, compared to DPO, DPO has achieved a performance improvement on IFEval (S-Acc, L-Acc) with Qwen2-7B as the base model. However, there are also cases of decline, such as comparing DPO with DPO on CFBench (CSR, PSR), which indicates that it is not the case that more tokens always lead to better performance. At the same time, although consuming the same number of tokens, SFT and DPO still have certain gaps compared to the proposed IOPO, which confirms that the performance improvement of our IOPO does not primarily come from using more tokens, but rather from better constraint-aware modeling of input-output preferences.
@@ 7 Conclusion
In Trace, the evaluation set has undergone strict manual verification but the training set has not performed this process considering the cost. Although the models trained on the training set have achieved the significant improvements, we believe that if we can further improve the quality of the training set, it will lead to better model performance on effectiveness and generalization of following complex instructions.
Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments and constructive suggestions. We sincerely thank members of the ConvAI Team in Tongyi Lab for their valuable feedback and discussions.
@@ Appendix B Prompt
B.1 Constraint Expansion Prompt
[Task Prompt]
@@ B.1 Constraint Expansion Prompt
B.2 Instruction Structuring Prompt
[Task Prompt]
@@ B.2 Instruction Structuring Prompt
{Instruction}
@@
B.3 Judge Completeness Prompt
[Task Prompt]
[Requirements]
The given instruction consists of three parts: <Task Description, Constraints, Input>, where Input can be NULL. You can refer to the examples given in Example, but you should not directly copy the examples.
[Example]
{Examples}
[Input]
@@ B.3 Judge Completeness Prompt
B.4 Judge Redundancy Prompt
[Task Prompt]
You are the redundancy detector for instructions, capable of determining whether given instructions are redundant.
[Requirements]
The given instruction consists of <Task Description, Constraints, Input>, where Input can be NULL; You can refer to the Examples provided, but you should not directly copy the examples.
[Example]
{Examples}
[Input]
@@ B.4 Judge Redundancy Prompt
{Instruction}
@@
B.5 Response Evaluation Prompt
@@
5. It is necessary to strictly follow the format in the [Example] for generation, the Fine Grained Score format is Json, and Constraints Overall Score format is List.
@@
[Example]
{Examples}
[Input]
@@ Appendix D IOPO Data Construction
Finally, we prompt Qwen2-72B-Instruct to check the rationality of the group pairs (<$x_{1},y_{1}$>, <$x_{2},y_{2}$>, <$x_{1},y_{2}$>, <$x_{2},y_{1}$>).
Model Method Trace IFEval CFBench ComplexBench
$\mathtt{IF}$-$\mathtt{S}$ $\mathtt{IF}$-$\mathtt{M}$ $\mathtt{S}$-$\mathtt{Acc}$ $\mathtt{L}$-$\mathtt{Acc}$ $\mathtt{CSR}$ $\mathtt{ISR}$ $\mathtt{PSR}$ $\mathtt{DRFR}$
Qwen2-7B Instruct 72.5 54.5 51.6 56.4 75.8 39.1 50.2 68.1
SFT 76.0 56.1 52.3 54.2 77.8 40.4 52.9 68.2
PPO 77.0 57.7 51.4 53.8 76.2 38.8 50.6 68.6
ORPO 77.9 61.7 53.1 56.9 79.7 45.9 57.0 69.1
SimPO 78.3 63.6 52.2 57.6 78.4 45.0 57.6 67.8
DPO 79.0 67.2 52.7 58.2 80.0 45.1 57.9 70.9
IOPO (Ours)Improv. 82.0↑3.0 68.9↑1.7 59.9↑7.2 63.6↑5.4 80.7↑0.7 47.0↑1.9 58.7↑0.8 72.6↑1.7
Llama3.1-8B Instruct 67.5 52.9 74.3 78.6 71.4 35.7 46.9 63.2
SFT 75.5 62.9 71.0 74.1 78.4 43.2 54.7 68.2
PPO 75.0 57.3 69.9 72.3 75.9 40.9 50.7 68.1
ORPO 77.0 63.1 72.3 77.3 79.4 46.6 57.2 69.6
SimPO 76.3 64.5 71.2 76.6 80.6 47.8 58.7 70.2
DPO 79.0 69.2 71.5 76.5 80.8 48.1 59.8 70.8
IOPO (Ours)Improv. 81.5↑2.5 70.7↑1.5 78.2↑6.7 81.0↑4.5 81.8↑1.0 49.9↑1.8 61.1↑1.3 71.8↑1.0
[…]
DeepSeek-R1 (671B) 84.1 73.8 83.5 89.4 86.5 62.1 71.9 78.7
Claude-3.5-Sonnet 87.9 76.3 86.7 90.5 87.1 62.6 72.3 82.6
Table 7: Performance on Trace, IFEval, CFBench, and ComplexBench.
Appendix E IOPO Variant
In fact, when implementing IOPO in the business applications, we can adopt the following simplification by omitting $\Pi_{2}(\pi_{\theta})$ in Eq.9 for less memory usage:
$\mathcal{L}_{\rm IOPO^{*}}(\pi_{\theta})$ $=-\mathbb{E}_{i\sim D}\,\bigg{\{}{\rm log}\bigg{[}\sigma\bigg{(}2\beta\log\frac{\pi_{\theta}(y_{1}|x_{1})}{\pi_{\rm ref}(y_{1}|x_{1})}$ (11)
$-\beta\log\frac{\pi_{\theta}(y_{2}|x_{1})}{\pi_{\rm ref}(y_{2}|x_{1})}-\beta\log\frac{\pi_{\theta}(y_{1}|x_{2})}{\pi_{\rm ref}(y_{1}|x_{2})}\bigg{)}\bigg{]}\bigg{\}}$
$i$ $=<x_{1},y_{1},x_{2},y_{2}>$
After simplification, the memory usage is approximately 1.5 times that of DPO, and the performance achieves average increase of 2.03% compared to DPO, decreases by 0.48% compared to before IOPO simplification.
Appendix F Compared to Mainstream LLMs
As shown in Table 7, we evaluate some mainstream closed-sourced/SOTA models, including OpenAI o1 Jaech et al. (2024), DeepSeek-R1 Guo et al. (2025), and Claude-3.5-Sonnet Anthropic (2024), and observe that: (1) Our benchmark Trace provides the consistent rank (Claude-3.5-Sonnet > o1 > DeepSeek-R1) as existing benchmarks IFEval, CFBench, and ComplexBench, confirming the data’s quality. (2) Larger closed-sourced/SOTA models have better performance than 7B/8B weaker models. However, the proposed algorithm IOPO significantly narrows the gap between closed-sourced/SOTA models and 7B/8B base weaker models, especially on in-domain Trace data. This inspires us to construct more diverse high-quality data and more advanced algorithm for enabling weaker models to achieve similar performance to larger-scale models.
Appendix G Derivation for $p(\mathcal{G}_{1}\succ\mathcal{G}_{2})$
@@
where the notations $p^{*}$ and $\pi^{*}$ are, respectively, equivalent to $p(\cdot)$ and $\pi_{\theta}$ in the main text.